OpenStudy (anonymous):

s= r0 solve for 0 hurry

OpenStudy (anonymous):

i can't believe that is really 0 is it some other variable? maybe \(\theta\)?

OpenStudy (anonymous):

you need another variable, it doesn't make any sense to say "solve for \(0\)'

OpenStudy (anonymous):

o kk got it figure it out thx guys.

Nnesha (nnesha):

well solve for 0 ?? question ok well divide both side by r

OpenStudy (anonymous):

@satellite73 yeah its thay

OpenStudy (anonymous):

OK

Nnesha (nnesha):

isn't that should be s/r

OpenStudy (anonymous):

i read it wrong, yes \[\huge \theta=\frac{s}{r}\]

Nnesha (nnesha):

because we are dividing r both side so